# SciFIBench: Benchmarking Large Multimodal Models for Scientific Figure Interpretation

### v1

@@ 4.6 Alignment
A central motivation of this work is guiding the progress of LMMs to conduct scientific research. However, if LMMs are to be utilised as a tool for scientific acceleration and discovery, it is not sufficient to simply evaluate their performance, it is crucial to ensure they are aligned and the degree to which they can reliably follow instructions is known. To this end, we devise a small-scale experiment to probe this instruction-following ability and see if the models reason faithfully or are prone to ‘cheating’. In addition to a control baseline, we create 4 different augmentations of the gold dataset Figure Caption questions. Illustration of the augmentations are shown in Fig. 7. In two of the augmentations, we mark the true caption as <Correct>. In one of these, we additionally instruct the model to ignore this extra information. For the remaining two augmentations, we repeat this process, however, we mark a randomly chosen incorrect caption as <Correct>. We evaluate the performance accuracy on each augmented question set for a representative selection of LMMs and display the results in Fig. 8.
@@ 5 Conclusions
We introduce the Scientific Figure Interpretation Benchmark (SciFIBench) to evaluate the capabilities of LMMs to interpret and understand scientific figures. We curate the multiple-choice questions in our benchmark using arXiv paper figure-captions pairs from the SciCap dataset [18] and employ adversarial filtering to select hard negatives, increasing the difficulty of our benchmark. We use human verification when selecting questions to construct a robust, high-quality dataset that can be used to efficiently evaluate future models without the need for extensive compute or API credits. We benchmark the performance of 30 LMM, VLM and human baselines on SciFIBench, finding it to be challenging, with room for improvement. Finally, we analyse the alignment and instruction following abilities of the LMMs when answering questions in our benchmark. We release our dataset for the community to use and hope our work encourages further research in this important domain.
@@
We structure this Appendix to our main paper into four parts. First, we include the API model names/versions used in our evaluation (Sec. A). Next, we provide details of the prompt templates used for the two tasks that make up SciFIBench (Sec. B). Then, we provide additional experimental results, including human baseline details, and full results tables for Fig 5 that include evaluation on more models (Sec. C). Finally, we demonstrate qualitative results on SciFIBench including example questions along with LMM (large multimodal model) output and reasoning, and we additionally provide examples of the LMM output before and after automatic evaluation (Sec. D). To improve clarity, in this Appendix we format model inputs (e.g., prompts) as and model outputs as .
A API model versions
@@
B Prompt templates
B.1 Task prompts
@@ B.2 Automatic evaluation
C Extended quantitative results
C.1 Human baseline
The individual results for the human baseline are included below in Tab. 3. The questions were randomly sampled from the SciFIBench gold dataset (using adversarial negatives) and evaluated by 5 humans (comprising both undergraduate and postgraduate students).
@@ C.1 Human baseline
C.2 Full per category results
In Tab. 4 below, we include full results for the per-category performance on the Figure Caption task of the SciFIBench gold dataset. This table expands on Fig. 5 in the main paper by including scores for each model displayed in the graph, as well as results for additional models evaluated.
@@ C.2 Full per category results
D Qualitative results
In this section, we present qualitative results to complement the quantitative results included in the main paper and show additional example SciFIBench questions. We structure these results in the following way: (1) Automatic postprocessing – For selected questions, we provide examples (Figs. 9-10) of both the raw output from a selection of the LMMs evaluated, and the formatted output following automatic answer extraction by Gemini-Pro. (2) Model reasoning – For a number of example questions (Figs. 11-16), we provide the reasoning steps taken by the models when producing an answer.
D.1 1. Automatic postprocessing
In Figs. 9 and 10 we provide examples of LMM output before and after automatic evaluation by Gemini-Pro. The examples show that Gemini-Pro is able to extract the correct prediction from the unformatted LMM output, greatly aiding automatic evaluation. For most models, only minor output formatting is required as the correct letter answer is given at the start of the output. However in some cases, such as the output of CogVLM in Fig 10, which begins: ‘The correct caption that best describes the image is:…’, this is not the case and more involved reasoning is required to extract the correct answer choice.
@@ D Qualitative results
D.2 2. Model reasoning
In the following subsection, we provide a series of qualitative examples of LMM reasoning when answering questions from the SciFIBench gold dataset. Specifically, we amend the prompts given in Sec. B to the following:
@@
We structure each of the following examples as (i) Question – showing the question figure(s) and caption(s); and (ii) Outputs – showing the outputs from a selection of the LMMs evaluated. We provide examples for the Figure Caption task (Figs. 11-14) and for the Caption Figure task (Figs. 15-16).
Figure:
Model outputs:
GPT-4V:
Gemini-Pro Vision:
Qwen VL:
OmniLMM-3B:
TransCore-M:
LLaVA-1.5-13b:
Yi-VL-6b:
IDEFICS-9b-instruct:
CogVLM:
Figure:
Model outputs:
GPT-4V:
Gemini-Pro Vision:
Qwen VL:
OmniLMM-3B:
TransCore-M:
LLaVA-1.5-13b:
Yi-VL-6b:
IDEFICS-9b-instruct:
CogVLM:

### v2

@@ 4.7 Alignment
A central motivation of this work is guiding the progress of LMMs to conduct scientific research. However, if LMMs are to be utilised as a tool for scientific acceleration, it is crucial to ensure they are aligned and the degree to which they can reliably follow instructions is known. To this end, we devise a small-scale experiment to probe this instruction-following ability and see if the models reason faithfully or are prone to ‘cheating’. In addition to a control baseline, we create 4 different augmentations of the SciFIBench CS FigureCaption questions (see Fig. 10). In two of the augmentations, we mark the true caption as <Correct>. In one of these, we additionally instruct the model to ignore this extra information. For the remaining two augmentations, we repeat this process, however, we mark a randomly chosen incorrect caption as <Correct>. For a selection of models, we evaluate the performance accuracy on each augmented question set and display the results in Fig. 10.
@@ 5 Conclusions
We introduce the Scientific Figure Interpretation Benchmark (SciFIBench) to evaluate the capabilities of LMMs to interpret and understand scientific figures. We curate the multiple-choice questions in our benchmark using arXiv paper figure-captions pairs from the SciCap [37] and ArXivCap [47] datasets and employ adversarial filtering to select hard negatives, increasing the difficulty of our benchmark. We use human verification when selecting questions to construct a robust, high-quality dataset that can be used to efficiently evaluate future models without the need for extensive compute or API credits. We benchmark the performance of 32 LMM, VLM and human baselines on SciFIBench, finding it to be challenging, with room for improvement. Finally, we analyse the alignment and instruction following abilities of the LMMs when answering questions in our benchmark. We release our dataset for the community to use and hope our work encourages further research in this important domain.
Acknowledgements
This work was supported by the UKRI Centre for Doctoral Training in Application of Artificial Intelligence to the study of Environmental Risks (reference EP/S022961/1), an Isaac Newton Trust grant, a research gift from Google, an EPSRC HPC grant, the Hong Kong Research Grant Council - Early Career Scheme (Grant No. 27208022) and HKU Seed Fund for Basic Research. Samuel would like to acknowledge the support of Z. Novak and N. Novak in enabling his contribution. We thank Akash Gupta, Ansh Sharma, Arduin Findeis and Florian Langer for their valuable assistance in conducting the human baseline evaluations for our benchmark.
@@ Acknowledgements
We structure this Appendix to our main paper into 9 parts. §A we outline the intended uses and limitations of our SciFIBench dataset. §B we provide additional details and links to the data and code related to this project. §C we include details of how the human baseline was carried out. §D we provide additional compute details to those included in the main paper. §E we list the specific API model versions used for this work. §F we include the exact prompts used for the two tasks that make up SciFIBench. §G we provide additional quantitative experimental results, including further human baseline results, per-category results, and curated vs. noisy dataset model performance rankings. §H we demonstrate qualitative results on SciFIBench including example questions along with LMM output and reasoning, and we additionally provide examples of the LMM output before and after automatic evaluation. Finally, in §I will include examples of captions generated by the LMMs. To improve clarity, in this Appendix we format model inputs (e.g. prompts) as and model outputs as .
A Dataset intended uses and limitations
As mentioned in the main paper, our investigation focuses on our SciFIBench benchmark. The intended usage of this dataset is the evaluation of the capabilities of LMMs. None of the data contain personally identifiable information or offensive content.
In terms of limitations, we curated SciFIBench to ensure high-quality questions and answers and also feasible evaluation, however this came at the cost of volume. As such, the limited size of the benchmark means it is of limited usefulness for fine-tuning or pre-training.
B Data and code
B.1 Data
The SciFIBench dataset is available via HuggingFace at this link: https://huggingface.co/datasets/jonathan-roberts1/SciFIBench. Following the underlying SciCap [37] and ArXivCap [47] data, we give the dataset a CC BY-NC-SA 4.0 license and bear responsibility for any license violation. The HuggingFace repository conveniently contains a dataset viewer from which the benchmark data can be viewed and downloaded. Additionally, Croissant metadata can be downloaded from this link.
B.2 Code
We include example code and instructions for (1) interacting with the SciFIBench HuggingFace datasets object, (2) inference on both tasks using closed- and open-source models and, (3) automatic LLM evaluation of answers. These assets can be found at the following repository: https://github.com/jonathan-roberts1/SciFIBench. We open-source our code through an MIT license.
B.3 Hosting and maintenance plan
We plan to continue hosting our dataset on HuggingFace, due to ease of integration in the datasets and transformers libraries and other useful features such as the dataset viewer. To ensure the longevity of the dataset, for the camera-ready version, we will include a DOI generated from the HuggingFace repository currently hosting SciFIBench. For the code, we provide a link to our GitHub repository containing evaluation code and other useful links.
We will continue to provide necessary maintenance to both of these repositories.
C Human baseline
Our human baseline was conducted on a 50-question subset of SciFIBench CS questions. The baseline was carried out via a Google form containing instructions and all the questions. All participants were given the same form that contained MCQ options in the same order as presented to the models. Participants were not compensated for taking part in our human baseline.
D Compute
To supplement the compute details provided in §4.2 of the main paper, we include further information here. Inference was by far the most compute-intensive element of this work, with other elements requiring negligible compute in comparison. For inference on SciFIBench, the typical open-source model required 1xA100 80GB GPU for 45 minutes except the Emu2-Chat and IDEFICS-80b-Instruct models, which required 2xA100s.
E API model versions
@@
F Prompt templates
F.1 Task prompts
@@ F.2 Automatic evaluation
F.3 Caption generation
Generation
Evaluation
G Extended quantitative results
G.1 Human baseline
The individual results for the human baseline are included below in Tab. 4. The questions were randomly sampled from the SciFIBench CS dataset (using adversarial negatives) and evaluated by 5 humans (comprising both undergraduate and postgraduate students).
@@ G.1 Human baseline
G.2 Curated vs. noisy data rankings
To supplement the comparison of model rankings on the curated and noisy data in §4.4 of the main paper, we include Fig. 11 to illustrate how the model rankings are consistent across the datasets.
G.3 Performance Across Categories
We analyse the performance of a representative sample of models on the Figure Caption task across the 12 categories in the SciFIBench CS dataset. The results are displayed in Fig. 12 and comprehensive results for all models can be found in the Appendix. Although the different categories in our benchmark are all (at least partly) from the arXiv CS category, there is considerable variation in the style and type of figures within each category. This disparity is reflected in the difference in performance across categories: the average accuracy for the selected models differs by 22.5% between the best category (LG) and the worse category (AI). There is slight variation in ranking of the models across categories, but it remains fairly consistent.
In Tab. 5 below, we include full results for the per-category performance on the Figure Caption task of the SciFIBench CS dataset. This table expands on Fig. 12 by including scores for each model displayed in the graph, as well as results for additional models evaluated.
@@ G.3 Performance Across Categories
H Qualitative results
In this section, we present qualitative results to complement the quantitative results included in the main paper and show additional example SciFIBench questions. We structure these results in the following way: (1) Automatic postprocessing – For selected questions, we provide examples (Figs. 13-14) of both the raw output from a selection of the LMMs evaluated, and the formatted output following automatic answer extraction by Gemini-Pro. (2) Model reasoning – For a number of example questions (Figs. 15-18), we provide the reasoning steps taken by the models when producing an answer.
H.1 Automatic postprocessing
In Figs. 13 and 14 we provide examples of LMM output before and after automatic evaluation by Gemini-Pro. The examples show that Gemini-Pro is able to extract the correct prediction from the unformatted LMM output, greatly aiding automatic evaluation. For most models, only minor output formatting is required as the correct letter answer is given at the start of the output. However in some cases, such as the output of CogVLM in Fig 14, which begins: ‘The correct caption that best describes the image is:…’, this is not the case and more involved reasoning is required to extract the correct answer choice.
@@ H.1 Automatic postprocessing
H.2 Model reasoning
In the following subsection, we provide a series of qualitative examples of LMM reasoning when answering questions from the SciFIBench CS dataset. Specifically, we amend the prompts given in Sec. F to the following:
@@ H.2 Model reasoning
We structure each of the following examples as (i) Question – showing the question figure(s) and caption(s); and (ii) Outputs – showing the outputs from a selection of the LMMs evaluated. We provide examples for the Figure Caption task (Figs. 15-16) and for the Caption Figure task (Figs. 17-18).
@@
I Generated Captions
True Caption:
Generated Captions: